The width of a rectangle is 6cm less than twice the length. If the area is to be 20cm^2, then what is its length? Show all work.
Answer by Cromlix(4381)

Hi there,
Make the length = x
Width = 2x - 6
Area = Length x Width
20 = x(2x - 6)
20 = 2x^2 - 6x
Rearrange:
2x^2 - 6x - 20 = 0
Divide by 2 for ease.
2(x^2 - 3x - 10)
Factorize
2(x + 2)(x - 5) = 0
x + 2 = 0
x = -2 (Disregard as -ve)
x - 5 = 0
x = 5
Length = 5 cm
Width = 4 cm
Hope this helps :-)